Hi, I am not sure how to set the equation up I guess, the answer I keep getting is 1518, answer key says it is wrong. The sales tax on a car that sold for $13,000 is $897. At the same rate, how much higher is the sales tax on a car that sells for $22,000?

Answer by ptfile(81)

Let $13,000 = base amount of the first car
$22,000 = base amount of the second car
$897 = Principal amount of the first car

The sales tax on a car that sold for $13,000 is $897.
Principal = (rate)(base)
897 = (r)(13,000)
897/13,000 = r
.069 or 6.9% = r
At the same rate, how much higher is the sales tax on a car that sells for $22,000?
Principal = (rate)(base)
P = (22,000)(.069)
p = 1518
Principal of second car - Principal of first car = difference of sales tax
1518 - 897 = $621
The sales tax on the second car is $621 higher than the first car.